# Interactive Program Synthesis for Modeling Collaborative Physical Activities from Narrated Demonstrations

### v1

@@ 4.2.1. Modeling Behavior as a Finite State Machine
The objective of program synthesis in our system is to model the user’s behavior from narrated demonstrations. We formally represent the behavior as a hierarchical, interrupt-driven finite state machine (FSM), a standard abstraction in program synthesis that compactly encodes contingencies via conditional branching. In our setting, states (nodes) denote actions, while directed edges denote preconditions—Boolean predicates over continuous variables (e.g. positions, orientations) of the physical world—that gate transitions. This abstraction allows the model to capture both discrete sequencing and coordinated behavior conditioned on the physical world state.
@@ 4.2.2. Library of APIs
Figure 4. LVLM output Scenic program for running example. Italicized identifiers (precondition1, etc) are learned constraints.:
4.3. Program synthesis
@@ 4.3.1. Data Processing
We overlay visual cues onto the RGB videos of narrated demonstrations. These cues include a grid to visualize object positions relative to the workspace and each other, markers for user annotations (e.g., pointing to where a player should move to create space for a pass), object labels, and arrows to indicate orientation. Such overlays support the LVLM’s visual and physical reasoning. Videos are then subsampled as a sequence of images at a uniform rate according to the LVLM’s token context length.
The transcribed narration is grounded with annotations and logged events from the demonstrations, with additional text injected into the transcript to clarify verbal references, establish action order, and enrich the description. For instance, in our running example, the transcript becomes: “You should move to either side of the field [user marked coordinate $(x,y)$] to open up for a pass. [teammate passed to user] Then shoot if a clear line is available [user shot toward the goal]; otherwise, pass back to your teammate.” Grounding is achieved by aligning timestamps of word tokens with logged events.
@@ 4.3.2. Program Generation
The grounded multimodal streams—comprising the annotated transcript and videos with overlaid visual cues—are provided as input to the LVLM together with the library of APIs described in Section 4.2.2. A system prompt guides the model to synthesize a Scenic program that encodes the demonstrated behavior utilizing the available APIs. In doing so, the LVLM need to account for both variations and overlap across narrated demonstrations, producing a program that generalizes the user’s intent while adhering to the FSM-based modeling scheme.
Adding Temporally Aligned Narration to System Demonstration. Demonstrations alone can be ambiguous, as users may infer multiple, conflicting interpretations of the system’s intent. To address this, we introduce a Speak() API that allows the program to narrate its intent during execution. This narration occurs in real-time, synchronized with the corresponding actions, enhancing interpretability. Prior work in interactive program synthesis has explored natural language explanations through inline annotations (Angert et al., 2023) or step-by-step summaries (Liu et al., 2023a). Building on this, our system prompts the LVLM during synthesis to insert line-by-line rationales using Speak(text) calls, which encode the reasoning behind each local action. At runtime, each Speak statement briefly pauses the simulation, plays the narration via a text-to-speech model, and then resumes execution. Because these narrations are embedded directly within the program’s structure, they are temporally aligned with the behavior being demonstrated, enabling users to understand not just what the system is doing, but why it is doing it, in the moment. Returning to our running example, the synthesized program modeling the user behavior with narration is shown in Fig. 4(left).
@@ 4.4.1. Decision Flow
Prior work shows that symbolic representations of program structure help non-programmers inspect logic and make targeted edits (Xie et al., 2024; Yen et al., 2024; Jiang et al., 2022; Kazemitabaar et al., 2024). We adopt this visualization approach by displaying the program’s FSM. For the program in Fig. 4, the user is shown the corresponding decision flow chart in Fig. 5(a). Nodes and edges include descriptions that users can reveal by clicking, which convey the rationale and underlying code logic without exposing the code itself. Users can annotate selected nodes and edges while providing verbal feedback, enabling direct edits to the program’s structure. The user might look at the decision flow chart in Fig. 5(a) and annotate the node for MoveTo and explain that ”You don’t need to move to the goal. Either make the shot or pass back.” Otherwise, if everything was fine with the decision flow chart, the user could just verbally state that ”no modifications are needed.”
@@ 4.4.2. Program Execution
The program runs as a Unity MR simulation, producing an embodied, narrated execution of the learned behavior. It controls an avatar—a proxy for the user—across variations of the environment encountered during teaching. As the behavior unfolds, the rationale and actions are narrated through the Speak actions defined in the code as seen in Fig. 4 (left). Users can pause at any time, point and click to annotate the environment, provide verbal feedback, and then resume—supporting multiple corrections within a single run. As in data processing for program synthesis, we capture both video and a grounded transcript of the user’s feedback, with interleaved annotations. Incorporating the Speak() outputs into the transcript further helps the LVLM localize feedback to specific points in the program logic, using the narration as a debug print statements. After looking at the program execution in Fig. 5(b) the user might pause at the frame shown and verbally explain that ”You should move further to the side of the field, around where I marked, to lure the opponent away [user annotated (x,y)(x,y)].”
@@ 4.4.3. Program Repair
We then provide the multimodal stream of feedback data including videos, annotations and grounded transcripts to the LVLM together with the candidate program to amend and the user’s initial narrated demonstration. The original prompt for program synthesis is augmented to instruct the LVLM on how to utilize feedback data and make localized edits.
Figure 5. User-in-the-loop feedback is supported through two complementary modes: (a) Decision Flow, which presents a finite state machine derived from the synthesized code, enabling structured inspection and annotation of program logic. Users are able to click and highlight edges and nodes by using ”Annotate.” When edges and nodes are highlighted yellow through annotation, our system will know which edgers and nodes need to be fixed. (b) System Execution, which replays the synthesized behavior under varied conditions with optional narrated Speak() actions, allowing users to validate fidelity in practice and provide situated corrections. Users can provide annotations in the environment by clicking on the field, leaving behind an X mark. Together, these modes balance interpretability and experiential validation, ensuring programs remain both logically consistent and behaviorally faithful.: :
@@
5. User Study
5.1. Participants
We recruited 20 participants via email to our institution’s intramural, club, and varsity soccer groups. Inclusion required competitive soccer experience (at least competitive intramural, club, or varsity at the high-school or college level). We excluded individuals who only played recreational leagues, had a history of seizures, visual abnormalities, susceptibility to motion sickness or nausea, scalp skin irritation, or used wearable medical devices, consistent with the MR headset’s safety guidance. Each participant received a \$30 gift card.
@@ 20 Semi-Professional 0 Male
5.3. Study Design
We conducted a within subjects study where participants evaluated the system’s learning before and after they provided corrective feedback. We chose soccer as an example domain for complex collaborative physical activities as it involves highly dynamic coordination with multiple players. Three different soccer scenarios programs were prepared by the study personnels, where the behaviors of the environment agents (e.g. teammates, opponents) were pre-scripted. These three scenarios are derived and adapted from the England Football Learning (England Football Learning, 2025), a popular website for youth soccer coaches regarding soccer drills. Each participant embodied the coach avatar in MR to provide narrated demonstrations. The soccer scenarios were assigned to participants in a round-robin manner.
5.4. Study Procedure
Each participant completed a 60-minute one-on-one session. After watching a tutorial video (covering how to teach in AR, inspect, and correct system learning), participants had a hands-on tutorial in MR to familiarize the buttons on the controllers. Then, they watched a video informing them of their assigned scenario and the pre-scripted behaviors of the environment agents (e.g. teammates, opponents), and reviewed the rubric that would be used for evaluation of the learned programs later to ensure that their teaching covers the rubric. They then recorded two narrated demonstrations of the same tactical goal (re-recording allowed). The system synthesized a program and displayed a decision-flow diagram; if issues were identified, participants were allowed to provide at most one corrective-feedback and the LLM repaired the program. Next, the system showed three narrated demos of its learned behavior; if problems remained, a fourth demo was shown and participants were allowed to provide at most one feedback, prompting another repair. The restrictions on the number of user feedback was enforced due to limited study time. Participants evaluated the program pre-feedback and post-feedback using the rubric. After viewing both versions’ decision-flow diagrams, they sketched their ground-truth decision flow and compared it with the system’s baseline and post-repair diagrams. Sessions concluded with a brief user experience interview (refer to tutorial videos and rubrics in the Supplement).
@@ 6.1. Metrics for Quantifying the Alignment of Learned Program Behaviors with User Intent
Completeness. This metric complements correctness. The program executions that users evaluated may not cover all branching code paths created by different conditional statements. Thus, completeness quantifies the alignment of the overall program structure (visualized as decision flow diagram) with the user intent. Recall, in the study, each user sketched their ground truth decision flow and compared it with the system’s 5.4. Let $G_{sys}=(V_{sys},E_{sys})$ denote the set of nodes and edges in the system-generated decision flow, and $G_{gt}=(V_{gt},E_{gt})$ denote the ground-truth’s. Completeness is calculated as the percentage of ground-truth elements (nodes and edges) that also appear in the system’s flow:
@@ 6.2. Statistical Analysis Method
Wilcoxon signed-rank test (Woolson, 2007) was used to evaluate to conduct one-sided hypothesis test on whether alignment improved post-feedback within subjects. Wilcoxon test was used instead of paired t-test because the resulting distribution of metric values were non-Gaussian. The scipy.stats.wilcoxon function from the SciPy library was used.
@@ 6.3. User Experience on Teaching via Embodied Narrated Demonstrations (RQ1)
95% of the users agreed that it was easy to teach soccer tactics via narrated demonstrations (Fig. 6). However, users also pointed out some limitations of the teaching interface, where two themes emerged. First, 8 out of 20 participants raised issues regarding the pre-scripted environment agents’ behaviors not aligning with their expectations to provide realistic teaching. The pointed out unrealistic movement (linear/stiff paths, stationary or slow defenders) and timing/latency that diverged from typical live play. For example, one noted “Players can only run in linear paths… addition of curved paths would be more realistic” (P2). Another shared “I didn’t feel that I was actually under pressure… unrealistic how slow [defenders are] moving” (P15). The other theme related to the lack of real-time feedback from the system to the participants while they were teaching. A few participants wanted two‑way feedback/acknowledgment from avatars during the teaching phase. “There’s no feedback from the [VR] players… felt a little off compared to real life.”(P13) “It felt a little bit like I was talking to the void… [I wanted] some way of having… feedback.” (P18)

### v2

@@ 4.2.1. Modeling Behavior as a Finite State Machine
The objective of program generation in our system is to model the user’s behavior from narrated demonstrations. We formally represent the behavior as a hierarchical, interrupt-driven finite state machine (FSM), a standard abstraction in program generation that compactly encodes contingencies via conditional branching. In our setting, states (nodes) denote actions, while directed edges denote preconditions—Boolean predicates over continuous variables (e.g. positions, orientations) of the physical world—that gate transitions. This abstraction allows the model to capture both discrete sequencing and coordinated behavior conditioned on the physical world state.
@@ 4.2.2. Library of APIs
4.3. Program Generation
@@ 4.3.1. Data Processing
We process three types of data collected during the teaching phase: (1) RGB video captured from a virtual camera, (2) timestamped action annotations triggered via MR controllers, and (3) narrated descriptions transcribed as timestamped tokens. The video is treated as a sequence of timestamped image frames. Inputting all frames directly to the LVLM could exceed its input token budget. In such a case, we sub-sample the sequence (e.g., selecting a frame every second) to preserve the temporal structure of the demonstration while keeping the number of images within the LVLM’s context limits.
We then ground the timestamped actions into the transcribed text so that the resulting transcript encodes both narration and user actions in a temporally synchronized form. For instance, in our running example, the grounded transcript becomes: “You should move to either side of the field [user marked coordinate (12.3, 2.45)] to open up for a pass. [Teammate1 passed to user] Then shoot if a clear line is available [user shoots toward the goal]; otherwise, pass back to your teammate.” Grounding is achieved by aligning timestamps of word tokens with logged events.
Figure 4. A snippet of a Scenic program modeling the user’s behavior when interacting with a teammate and an opponent in the running soccer example. The actions, constraints, and preconditions are expressed using a provided API library that defines the action space and physical constraints described in Sec. 4.2.2.:
@@ 4.3.2. Program Generation
The LVLM takes the following as input: (1) pairs of sub-sampled images and their corresponding grounded transcripts, (2) textual annotations of the available APIs, (3) documentation of the syntax and semantics of the target programming language, and (4) a system prompt. The system prompt provides a few example programs and instructs the LVLM to generate a Scenic program that models the user’s instructions using the provided APIs. For concrete details of these inputs, please refer to Supplement A.7. Through this query, the LVLM generates a full executable Scenic program, instantiating agents and objects and assigning initial conditions and behaviors to agents.
Adding Temporally Aligned Narration to System Demonstration. Demonstrations alone can be ambiguous, as users may infer multiple, conflicting interpretations of the system’s intent. To address this, we introduce a Speak() API that allows the program to narrate its intent during execution. This narration occurs in real-time, synchronized with the corresponding actions, enhancing interpretability. Prior work in interactive program generation has explored natural language explanations through inline annotations (angert2023spellburst) or step-by-step summaries (liu2023whatitwantstosay). Building on this, our system prompts the LVLM during generation to insert line-by-line rationales using Speak(text) calls, which encode the reasoning behind each local action. At runtime, each Speak statement briefly pauses the simulation, plays the narration via a text-to-speech model, and then resumes execution. Because these narrations are embedded directly within the program’s structure, they are temporally aligned with the behavior being demonstrated, enabling users to understand not just what the system is doing, but why it is doing it, in the moment. Returning to our running example, the generated program modeling the user behavior with narration is shown in Fig. 4(left).
Figure 5. User-in-the-loop feedback is supported through two complementary modes: (a) Decision Flow, which presents a finite state machine derived from the generated code, enabling structured inspection and annotation of program logic. Users are able to click and highlight edges and nodes by using ”Annotate.” When edges and nodes are highlighted yellow through annotation, our system will know which edgers and nodes need to be fixed. (b) System Execution, which replays the modeled behavior under varied conditions with optional narrated Speak() actions, allowing users to validate fidelity in practice and provide situated corrections. Users can provide annotations in the environment by clicking on the field, leaving behind an X mark. Together, these modes balance interpretability and experiential validation, ensuring programs remain both logically consistent and behaviorally faithful.: :
@@ 4.4.1. Decision Flow
Prior work shows that symbolic representations of program structure help non-programmers inspect logic and make targeted edits (xie2024waitgpt; yen2024coladder; jiang2022genline; kazemitabaar2024interactive). We adopt this visualization approach by displaying the program’s FSM. For the program in Fig. 4, the user is shown the corresponding decision flow chart in Fig. 5(a). Nodes and edges include descriptions that users can reveal by clicking, which convey the rationale and underlying code logic without exposing the code itself. Users can annotate selected nodes and edges while providing verbal feedback, enabling direct edits to the program’s structure. The user might look at the decision flow chart in Fig. 5(a) and annotate the node for MoveTo and explain that ”You don’t need to move to the goal. Either make the shot or pass back.” Otherwise, if everything was fine with the decision flow chart, the user could just verbally state that ”no modifications are needed.”
@@ 4.4.2. Program Execution
The program runs as a Unity MR simulation, producing an embodied, narrated execution of the learned behavior. It controls an avatar—a proxy for the user—across variations of the environment encountered during teaching. As the behavior unfolds, the rationale and actions are narrated through the Speak actions defined in the code as seen in Fig. 4 (left). Users can pause at any time, point and click to annotate the environment, provide verbal feedback, and then resume—supporting multiple corrections within a single run. As in data processing for program generation, we capture both video and a grounded transcript of the user’s feedback, with interleaved annotations. Incorporating the Speak() outputs into the transcript further helps the LVLM localize feedback to specific points in the program logic, using the narration as a debug print statements. After looking at the program execution in Fig. 5(b) the user might pause at the frame shown and verbally explain that ”You should move further to the side of the field, around where I marked, to lure the opponent away [user annotated (x,y)(x,y)].”
@@ 4.4.3. Program Repair
We then provide the multimodal stream of feedback data including videos, annotations and grounded transcripts to the LVLM together with the candidate program to amend and the user’s initial narrated demonstration. The original prompt for program generation is augmented to instruct the LVLM on how to utilize feedback data and make localized edits.
5. Generalizability of Our Approach
Figure 6. Design space of collaborative physical activities along seven dimensions. Soccer-ball and robot emojis mark values instantiated in soccer and factory examples ( Fig. 5 and 7), respectively; yellow cells indicate dimension values not supported by our current system, while green cells indicate dimension values that are supported.
Overall scope of activities. Our approach can model a broad class of collaborative physical activities in which an individual agent’s behavior is represented as a high-level policy over physical states and actions. It reasons about how a focal agent should react to or coordinate with others by planning and sequencing such actions (e.g., “dribble,” “grasp an object,” “move into open space”). However, it cannot model the underlying low-level motor skills that implement these actions, which require fine-grained coordination among joints with many degrees of freedom. It is unclear how to expose this kind of motor-domain knowledge as reusable APIs; for example, specifying how finger joints should coordinate in 3D to grasp objects with diverse shapes, sizes, and tactile properties. Without APIs that encode this level of detail, it is difficult to learn such coordination from only a few narrated demonstrations, and learning such fine-grained motor skills from even large datasets remains an open challenge in robotics. Furthermore, our approach is not designed to model activities that aim to teach a team of collaborators as a single unit. For user inspection (Sec. 4.3.2), the system generates narrated demonstrations that explain the rationale behind the actions of a single agent; it is not designed to model or explain collective coordination strategies across multiple agents. In practice, this means our approach is applicable for modeling high-level policies of individual agents acting within collaborative activities, given a predefined low-level action space, but not for learning the underlying motor skills or synthesizing team-level control policies.
Dimensions of collaborative activities. To more precisely characterize the class of activities our system targets within this scope, we adopt seven dimensions of collaborative activities grounded in prior work (Table 11). Goal Alignment (Deutsch1949; JohnsonJohnson2005) captures whether the task goal is shared among collaborators (e.g., manufacturing) or conflicting (e.g., one-on-one boxing). Agent Composition (DeSantis2008; OnnaschRoesler2021) distinguishes collaboration among humans or robots. Instruction Scope (MaloneCrowston1994; Gervasi2020) specifies whether the task targets an individual collaborating with others or a team working collectively. Role Structure (Belbin1981; Belbin1993) characterizes whether collaborators’ roles are identical (e.g., synchronized dancing) or heterogeneous (e.g., soccer). Coordination Level (DeSantis2008; MaloneCrowston1994) indicates whether the collaboration requires coordination at the planning level or the control level. Here, control refers to low-level actions that require coordinating multiple body joints with high degrees of freedom (e.g., grasping or throwing a ball), while planning sequences these low-level actions to achieve a task. Control Authority (Parasuraman2000; Sheridan1992) refers to who controls collaborators’ behaviors, ranging from pre-determined policies (e.g., pre-determined strategies of soccer opponents) to real-time user control. Finally, Communication Modality (OnnaschRoesler2021; Gervasi2020) specifies whether interaction relies on linguistic or gestural communication among agents (e.g., raising a hand to signal another agent). Combinations of values within each dimension are also possible but are omitted from the table for brevity.
Figure 7. User feedback on decision flow and system execution in a factory manufacturing setting where a human and a robot collaborates to package a box.
Instantiated examples: soccer and factory. Within this design space, we instantiated our system in two concrete domains. The running soccer example in Sec. 4 models multi-agent tactics in which a focal player coordinates with teammates and opponents at the level of high-level decisions (e.g., luring defenders, creating passing lanes). To further demonstrate generalizability, as illustrated in Fig. 7, we instantiated our system architecture in a factory manufacturing setting in which a mobile robot collaborates with a human worker to package a box. In this scenario, the user authors the interaction by instantiating the environment (placing the human, robot, and box) and specifying initial conditions and behaviors for each agent. For example, the user can define that the human begins by raising a hand to gesture to the robot, and that the robot should interpret this gesture as a command to fetch a box, deliver it to the human, and enable the human to complete the packaging task. Following the procedures described in Sec. 4, the user then teaches, inspects, and corrects the programs modeling this collaboration. We present an existence-proof with three participants, all of whom were able to author and iteratively refine collaborative activities, yielding programs that accurately modeled their desired human–robot interactions. The results are included in Supplement B.
In Table 11, the cells highlighted in green and marked with soccer-ball or robot emojis indicate which dimension values are instantiated in the soccer and factory examples, respectively. Cells where both values of a dimension are highlighted in green indicate that our approach can model combinations of values, which are omitted in the table.
Unsupported and untested regions of the space. Cells highlighted in yellow denote dimension values that our system cannot currently model. Cells that are not colored correspond to dimension values that we did not empirically evaluate in this paper but for which we expect our approach to generalize. For example, homogeneous role structures such as synchronized dancing could be modeled by assigning the same learned behavior to all agents in the activity. Likewise, linguistic communication among agents could be supported by using large language models to parse utterances and trigger appropriate coordinated actions.
6. User Study
6.1. Participants
The protocol for our study was approved by our institution’s IRB. We recruited 20 participants via email to our institution’s intramural, club, and varsity soccer groups. Inclusion criteria required competitive soccer experience (at least competitive intramural, club, or varsity at the high-school or college level). We excluded individuals who only played recreational leagues, had a history of seizures, visual abnormalities, susceptibility to motion sickness or nausea, scalp skin irritation, or used wearable medical devices, consistent with the MR headset’s safety guidance. The signed informed consent form was obtained for each subject via email or in person prior to the study. Each participant received a \$30 gift card.
@@ 6.2. Experiment Setup
6.3. Study Design
We conducted a within subjects study where participants evaluated the system’s learning before and after they provided corrective feedback. We chose soccer as an example domain for complex collaborative physical activities as it involves highly dynamic coordination with multiple players. Three different soccer scenarios (luring opponent, overlapping teammate, distributing pass) were prepared by the study personnels (refer to Supplement A.2), where the behaviors of the environment agents (e.g. teammates, opponents) were pre-scripted. These three scenarios are derived and adapted from the England Football Learning (englandfootball-sessions), a popular website for youth soccer coaches regarding soccer drills. Each participant embodied the coach avatar in MR to provide narrated demonstrations. The soccer scenarios were assigned to participants in a round-robin manner.
6.4. Study Procedure
Each participant completed a 60-minute one-on-one session. After watching a tutorial video (covering how to teach in AR, inspect, and correct system learning), participants had a hands-on tutorial in MR to familiarize the buttons on the controllers. Then, they watched a video informing them of their assigned scenario and the pre-scripted behaviors of the environment agents (e.g. teammates, opponents), and reviewed the rubric that would be used for evaluation of the learned programs later to ensure that their teaching covers the rubric. This rubric per scenario specified objective task requirements, such as scoring a goal and passing to a teammate (refer to Supplement A.3). They then recorded two narrated demonstrations of the same tactical goal (re-recording allowed). The system synthesized a program and displayed a decision-flow diagram; if issues were identified, participants were allowed to provide at most one corrective-feedback and the LLM repaired the program. Next, the system showed three narrated demos of its learned behavior; if problems remained, a fourth demo was shown and participants were allowed to provide at most one feedback, prompting another repair. The restrictions on the number of user feedback was enforced due to limited study time. Furthermore, we asked users to provide feedback that was consistent with their initial narrated demonstrations, even if they came up with better tactics over time, so that we could clearly measure improvement before and after feedback within the limited session. Participants evaluated the program pre-feedback and post-feedback using the rubric. After viewing both versions’ decision-flow diagrams, they sketched their ground-truth decision flow and compared it with the system’s baseline and post-repair diagrams. Sessions concluded with a brief user experience interview (refer to tutorial videos and rubrics in the Supplement).
@@ 7.1. Metrics for Quantifying the Alignment of Learned Program Behaviors with User Intent
Completeness. This metric complements correctness. The program executions that users evaluated may not cover all branching code paths created by different conditional statements. Thus, completeness quantifies the alignment of the overall program structure (visualized as decision flow diagram) with the user intent. Recall, in the study, each user sketched their ground truth decision flow and compared it with the system’s 6.4. Let $G_{sys}=(V_{sys},E_{sys})$ denote the set of nodes and edges in the system-generated decision flow, and $G_{gt}=(V_{gt},E_{gt})$ denote the ground-truth’s. Completeness is calculated as the percentage of ground-truth elements (nodes and edges) that also appear in the system’s flow:
@@ 7.2. Statistical Analysis Method
Wilcoxon signed-rank test (Woolson2007) was used to evaluate to conduct one-sided hypothesis test on whether alignment improved post-feedback within subjects. Wilcoxon test was used instead of paired t-test because participants were sampled from non-Gaussian distribution, with particular set of skills (e.g. playing in competitive or semi-professional soccer leagues). The scipy.stats.wilcoxon function from the SciPy library was used.
@@ 7.3. User Experience on Teaching via Embodied Narrated Demonstrations (RQ1)
Users authored programs modeling diversity of tactics per scenario (refer to Supplement A.6). 95% of the users agreed that it was easy to teach soccer tactics via narrated demonstrations (Fig. 8). However, users also pointed out some limitations of the teaching interface, where two themes emerged. First, 8 out of 20 participants raised issues regarding the pre-scripted environment agents’ behaviors not aligning with their expectations to provide realistic teaching. The pointed out unrealistic movement (linear/stiff paths, stationary or slow defenders) and timing/latency that diverged from typical live play. For example, one noted “Players can only run in linear paths… addition of curved paths would be more realistic” (P2). Another shared “I didn’t feel that I was actually under pressure… unrealistic how slow [defenders are] moving” (P15). The other theme related to the lack of real-time feedback from the system to the participants while they were teaching. A few participants wanted two‑way feedback/acknowledgment from avatars during the teaching phase. “There’s no feedback from the [VR] players… felt a little off compared to real life.”(P13) “It felt a little bit like I was talking to the void… [I wanted] some way of having… feedback.” (P18)